# Multi-Metric Preference Alignment for Generative Speech Restoration

### v1

@@ 1 Introduction
Speech restoration, which aims to recover high-quality speech from various degradations, is a fundamental task in audio processing (Liu et al. 2022; Zhang et al. 2024b). Recent advances in utilizing generative models for speech restoration, namely generative speech restoration (GenSR), have shown remarkable performance in tasks such as denoising, dereverberation, declipping, super-resolution, etc. (Zhang et al. 2025a; Yang et al. 2024b; Li, Wang, and Liu 2024; Wang et al. 2025c, 2024b; Yao et al. 2025a; Kang et al. 2025; Wang et al. 2024a; Liu et al. 2024b). Unlike traditional discriminative methods that minimize a distance metric to a clean reference, generative speech restoration models are trained to maximize the likelihood of clean speech given a degraded input, thus learning the underlying distribution of clean speech (Lemercier et al. 2023), enabling them to generate high-fidelity audio even from severely degraded inputs.
In parallel, post-training techniques, designed to align generative models with specific downstream objectives or human preferences, have become integral to the advancement of generative modeling in some domains. In natural language processing (Ouyang et al. 2022; Rafailov et al. 2023; Bai et al. 2022), text-to-image/video (Xu et al. 2023; Wallace et al. 2024; Liu et al. 2025), and text-to-speech/audio (Zhang et al. 2025b; Sun et al. 2025; Hussain et al. 2025; Tian et al. 2025; Yao et al. 2025b; Zhang et al. 2024a; Liao et al. 2024) synthesis, preference-based alignment methods have been instrumental in enhancing model quality, safety, and human alignment. However, despite its potential, the application of such post-training alignment to the domain of generative speech restoration remains largely under-explored.
@@
To address these challenges, we propose and investigate a multi-metric preference alignment strategy. We argue that a robust solution to reward hacking lies in the preference signal itself being multi-dimensional and holistic. To this end, we construct a new preference dataset, which we name GenSR-Pref. A sample is chosen as winner in GenSR-Pref only when it is unanimously judged superior by a complementary suite of metrics that assess four distinct aspects of quality: perceptual quality, signal-level fidelity, content alignment, and timbre preservation. This strict, unanimous criterion ensures that our model learns from a robust and holistic preference signal, effectively mitigating the risk of reward hacking.
Based on the GenSR-Pref dataset, we apply Direct Preference Optimization (DPO) (Rafailov et al. 2023) to align models across three major distinct generative paradigms: the sequential Autoregressive (AR) models, the iterative Masked Generative (MGM) models (also known as masked or discrete diffusion), and the continuous Flow-Matching (FM) models. Our approach yields significant and consistent improvements across all benchmarks even with only 3k pairs, confirmed by objective and subjective tests. Our key findings are threefold. First, our multi-metric signal is crucial for avoiding the reward hacking inherent in single-metric optimization. Second, we show that using ground-truth as a fixed winner causes model collapse, highlighting the need for learned, relative preferences over absolute targets. Third, models perform best with data from their own architecture, a principle we term in-paradigm alignment. We attribute this to unique “alignment directions” for each paradigm, where in-paradigm data provides a more direct optimization path. Finally, we demonstrate that our aligned models can act as powerful “data annotators”, generating high-quality pseudo-labels to train discriminative models in data-scarce scenarios like singing voice restoration, thus bridging the gap between generative and discriminative paradigms.
@@ 2.1 Generative Speech Restoration
The paradigm for speech restoration has recently shifted towards generative modeling, which can be broadly categorized by the underlying generation mechanism. One major stream leverages autoregressive language models to treat restoration as a sequence-to-sequence problem, as seen in SELM (Wang et al. 2024b), TSELM (Tang, Zeng, and Li 2024), GenSE (Yao et al. 2025a), and versatile multi-task systems like SpeechX (Wang et al. 2024a), UniAudio (Yang et al. 2024a), and LLaSE-G1 (Kang et al. 2025). A second family consists of masked generative models, or discrete diffusion models, which iteratively refine tokenized audio representations. This approach has been applied to general restoration (e.g., MaskSR (Li, Wang, and Liu 2024; Liu et al. 2024c), AnyEnhance (Zhang et al. 2025a)), speech enhancement (e.g., Genhancer (Yang et al. 2024b)), and target speaker extraction (e.g., Metis (Wang et al. 2025c)). A third direction employs generative models with continuous dynamics to learn mappings in a continuous space; this includes methods based on score-SDE (Welker, Richter, and Gerkmann 2022), latent diffusion (Liu et al. 2024b; Wang et al. 2025a), flow-matching (Wang et al. 2025d; Liu et al. 2024a; Ku et al. 2024), and schrödinger bridges (Li et al. 2025a; Jukić et al. 2024) for various restoration tasks. Despite achieving state-of-the-art results, these generative models are typically trained using likelihood-based objectives, which often misalign with human perceptual quality. Our work addresses this gap by introducing a preference-based post-training strategy applicable across these diverse generative paradigms.
@@ 2.2 Post-Training for Audio Alignment
Post-training alignment has proven effective across various generative domains, including text (Ouyang et al. 2022; Rafailov et al. 2023; Bai et al. 2022), vision (Xu et al. 2023; Wallace et al. 2024; Liu et al. 2025), speech (Zhang et al. 2025b; Sun et al. 2025; Hussain et al. 2025; Tian et al. 2025; Yao et al. 2025b; Zhang et al. 2024a), music (Cideron et al. 2024; Lei et al. 2025), and audio effects (Majumder et al. 2024; Liao et al. 2024). In the context of speech restoration, early efforts like MetricGAN (Fu et al. 2019) used adversarial training to optimize for metrics like PESQ. More recently, preference-based methods have emerged, such as aligning with a learned NISQA predictor via PPO (Kumar, Perrault, and Williamson 2025) or using DPO with a single metric like UTMOS (Li et al. 2025b). While single-metric alignment is effective, it risks reward hacking by ignoring that robust restoration is multi-faceted, requiring a balance of quality, fidelity, and timbre (Section 4.3). Grounded in this principle, we propose a multi-metric alignment strategy to pursue holistic improvement. We systematically show this approach yields more comprehensive gains across diverse generative paradigms while naturally mitigating reward hacking.
@@ 3.1 The GenSR-Pref Dataset: Curating a Holistic Preference Signal
Masked Generative Model (MGM). We use the pre-trained AnyEnhance (Zhang et al. 2025a) model as our MGM model. It predicts acoustic tokens based on the DAC codec (Kumar et al. 2024) from partially masked sequences, effectively capturing non-autoregressive and context-aware dependencies.
Autoregressive Model (AR). We train a new model, AR+Soundstorm, which follows a two-stage pipeline common in text-to-speech (TTS) domain. It first predicts semantic tokens of the Metis codec (Wang et al. 2025c) given noisy inputs, and a second-stage non-autoregressive model similar to Soundstorm (Borsos et al. 2023) then converts these tokens into acoustic tokens for waveform synthesis.
Flow-Matching Model (FM). We train Flow-SR, a model that learns to predict the velocity field of clean mel-spectrograms from noisy inputs via optimal transport interpolation based on the Diffusion Transformer (DiT) architecture (Peebles and Xie 2023). A pre-trained Vocos vocoder (Siuzdak 2023) is then used to synthesize the final waveform from the generated spectrograms. Details on model architectures and training procedures can be found in Appendix B.
@@
  • Perceptual Quality (NISQA) (Mittag et al. 2021): Assesses overall listening quality, including naturalness and the absence of annoying artifacts.
  • Signal-level Fidelity (DNSMOS) (Reddy, Gopal, and Cutler 2022): A composite metric evaluating signal distortion, background noise, and overall quality for a fine-grained assessment of signal integrity.
  • Content Alignment (SpeechBERTScore) (Saeki et al. 2024): Measures semantic similarity to the ground-truth transcription to ensure linguistic content is not altered during restoration.
@@ 3.2 GenSR Model Alignment via Direct Preference Optimization
In the second stage of our approach, we align the generative speech restoration models using the GenSR-Pref dataset. We employ Direct Preference Optimization (DPO) (Rafailov et al. 2023), a simple yet powerful technique that directly optimizes a policy to satisfy preferences without explicit reward modeling or reinforcement learning. The core idea of DPO is to re-parameterize the reward function in the standard RLHF objective using the optimal policy, which leads to a simple contrastive loss over preferred and dispreferred samples. In this section, we detail how we adapt and apply DPO to the three diverse generative paradigms used in our study: Autoregressive (AR), Masked Generative (MGM), and Flow-Matching (FM) models. Detailed derivations can be found in Appendix A.
@@
MGM models (Chang et al. 2022; Zhang et al. 2025a; Wang et al. 2025c; Ju et al. 2024; Wang et al. 2025b), also known as discrete diffusion models, learn to restore a clean speech sequence $y_{0}$ from a partially masked or corrupted version $y_{t}$, where $t\in[0,T]$ is the noise step parameter that controls the amount of corruption, and $y_{T}$ is the fully masked sequence. They are trained to predict the original tokens in the masked positions, optimizing a masked language modeling objective:
@@
where $m_{t}\in\{0,1\}^{T}$ is a binary mask indicating which tokens in the sequence are masked (1) or unmasked (0). The model learns to predict the original tokens in the masked positions, effectively learning to denoise the corrupted input.
DPO Derivation. Following the INTP framework (Zhang et al. 2025b), we extend DPO to this non-autoregressive setting. The core logic remains the same, but the policy now represents the conditional probability of the full clean sequence $y_{0}$ given the corrupted version $y_{t}$. The DPO loss for MGM is thus analogous to the AR case, contrasting the log-probabilities of the preferred and dispreferred sequences, conditioned on their respective masked inputs:
@@
Here, $y_{t}^{w}$ and $y_{t}^{l}$ are the masked versions of the winning ($y_{0}^{w}$) and losing ($y_{0}^{l}$) sequences from our preference dataset.
@@
FM models (Lipman et al. 2022) learn a continuous mapping from a simple prior distribution (e.g., Gaussian noise $y_{0}\sim\mathcal{N}(0,I)$) to a complex data distribution (e.g., clean speech $y_{1}$). They are trained to predict the velocity vector field $v_{\theta}(y_{t},t,x)$ that describes the probabilistic flow of data points along a linear interpolation path from the noise $y_{0}$ to the data $y_{1}$. The standard training objective minimizes the difference between the predicted velocity and the true velocity:
@@
DPO Derivation. To extend DPO to FM models, we follow recent advances in aligning diffusion-based generators (Wallace et al. 2024; Liu et al. 2025). Instead of modeling the full data likelihood along the entire generative trajectory, we use a tractable surrogate objective that operates on a single noise step. Specifically, we approximate the DPO objective using the model’s instantaneous prediction error at a sampled timestep, treating the L2 loss between predicted and true velocities as a proxy for negative log-likelihood. The resulting loss encourages the model to reduce velocity error for preferred samples ($y_{w}$) while increasing it for dispreferred ones ($y_{l}$), yielding the following expression:
@@ 4.1 Experimental Setup
Evaluation Benchmarks and Metrics. We conduct evaluations on two major kinds of benchmarks: (1) General speech restoration (GSR) that includes denoising, dereveration, declipping and super-resolution at the same time (Voicefixer-GSR, Librivox-GSR, CCMusic-GSR), (2) Speech Enhancement (SE) under noisy and reverberant conditions (DNS-No-Reverb and DNS-With-Reverb). For a comprehensive evaluation, we not only include the original unaligned models, but also involve state-of-the-art baselines LLaSE-G1 (Kang et al. 2025), GenSE (Yao et al. 2025a), MaskSR (Li, Wang, and Liu 2024), and FlowSE (Wang et al. 2025d), and a discriminative model Voicefixer (Liu et al. 2022). Performance is measured using a comprehensive suite of objective metrics: DNSMOS (SIG, BAK, OVRL) (Reddy, Gopal, and Cutler 2022) for signal fidelity and quality, NISQA (Mittag et al. 2021) for perceptual quality, SpeechBERTScore (SBERT) (Saeki et al. 2024) for content consistency, and speaker similarity (SIM). For subjective evaluation, we conduct A/B preference tests to assess human perception. Details on the evaluation can be found in Appendix C.
@@ 4.3 Ablation Studies and Analysis
The Role of Data Source in Alignment. Finally, we explore how the source of preference data affects alignment performance. Table 4 shows the results of aligning each model with preference pairs generated from itself (in-paradigm) versus from other paradigms (cross-paradigm). While models can often benefit from cross-paradigm data, we observe a clear trend: all three models achieve their optimal performance when aligned with their own in-paradigm preference data. To quantitatively investigate this phenomenon, we formally define the principle of in-paradigm alignment. We hypothesize that each architecture possesses a unique “preference direction”—an optimal path for improvement. To quantify this, we define a “preference vector” as the feature-space difference between winner and loser samples ($v_{\text{pref}}=f(y_{w})-f(y_{l})$) using pre-trained w2v-bert-2 model [^2] https://huggingface.co/facebook/w2v-bert-2.0 and compute their average pairwise cosine similarity. The results (Table 5) provide strong quantitative evidence for our hypothesis: the average in-paradigm similarity (diagonal values) is consistently higher than the cross-paradigm similarity (e.g., AR-AR similarity of 0.0075 vs. AR-MGM of 0.0018). This indicates that data from a model’s own paradigm provides a more consistent and targeted optimization signal.

### v2

@@ 1 Introduction
Speech restoration, which aims to recover high-quality speech from various degradations, is a fundamental task in audio processing (liu2022voicefixer; zhang24h_interspeech). Recent advances in utilizing generative models for speech restoration, namely generative speech restoration (GenSR), have shown remarkable performance in tasks such as denoising, dereverberation, declipping, super-resolution, etc. (zhang2025anyenhance; yang2024genhancer; li2024masksr; wang2025metis; wang2024selm; anonymous2025gense; kang2025llase; wang2024speechx; liu2024audiosr). Unlike traditional discriminative methods that minimize a distance metric to a clean reference, generative speech restoration models are trained to maximize the likelihood of clean speech given a degraded input, thus learning the underlying distribution of clean speech (lemercier2023analysing), enabling them to generate high-fidelity audio even from severely degraded inputs.
In parallel, post-training techniques, designed to align generative models with specific downstream objectives or human preferences, have become integral to the advancement of generative modeling in some domains. In natural language processing (ouyang2022instructgpt; rafailov2023direct; bai2022rlhf), text-to-image/video (xu2023imagereward; wallace2024diffusiondpo; liu2025videoalign), and text-to-speech/audio (zhang2025advancing; sun2025f5r; hussain2025koel; tian2025preference; yao2025fine; zhang2024speechalign; liao2024baton) synthesis, preference-based alignment methods have been instrumental in enhancing model quality, safety, and human alignment. However, despite its potential, the application of such post-training alignment to the domain of generative speech restoration remains largely under-explored.
@@
To address these challenges, we propose and investigate a multi-metric preference alignment strategy. We argue that a robust solution to reward hacking lies in the preference signal itself being multi-dimensional and holistic. To this end, we construct a new preference dataset, which we name GenSR-Pref. A sample is chosen as winner only when it is unanimously judged superior by a complementary suite of metrics that assess four distinct aspects of quality: perceptual quality, signal-level fidelity, content alignment, and timbre preservation. This strict, unanimous criterion ensures that our model learns from a robust and holistic preference signal, effectively mitigating the risk of reward hacking.
Based on the GenSR-Pref dataset, we apply Direct Preference Optimization (DPO) (rafailov2023direct) to align models across three major distinct generative paradigms: the sequential Autoregressive (AR) models, the iterative Masked Generative (MGM) models (also known as masked or discrete diffusion), and the continuous Flow-Matching (FM) models. Our approach yields significant and consistent improvements across all benchmarks even with only 3k pairs, confirmed by objective and subjective tests. Our key findings are threefold. First, our multi-metric signal is crucial for avoiding the reward hacking inherent in single-metric optimization. Second, we show that using ground-truth as a fixed winner causes model collapse, highlighting the need for learned, relative preferences over absolute targets. Third, models perform best with data from their own architecture, a principle we term in-paradigm alignment. We attribute this to unique “alignment directions” for each paradigm, where in-paradigm data provides a more direct optimization path. Finally, we demonstrate that our aligned models can act as powerful “data annotators”, generating high-quality pseudo-labels to train discriminative models in data-scarce scenarios like singing voice restoration, thus bridging the gap between generative and discriminative paradigms.
@@ 2.1 Generative Speech Restoration
The paradigm for speech restoration has recently shifted towards generative modeling, which can be broadly categorized by the underlying generation mechanism. One major stream leverages autoregressive language models to treat restoration as a sequence-to-sequence problem, as seen in SELM (wang2024selm), TSELM (tang2024tselm), GenSE (anonymous2025gense), and versatile multi-task systems like SpeechX (wang2024speechx), UniAudio (Yang2024UniAudio), and LLaSE-G1 (kang2025llase). A second family consists of masked generative models, or discrete diffusion models, which iteratively refine tokenized audio representations. This approach has been applied to general restoration (e.g., MaskSR (li2024masksr; liu2024joint), AnyEnhance (zhang2025anyenhance)), speech enhancement (e.g., Genhancer (yang2024genhancer)), and target speaker extraction (e.g., Metis (wang2025metis)). A third direction employs generative models with continuous dynamics to learn mappings in a continuous space; this includes methods based on score-SDE (welker2022sgmse), latent diffusion (liu2024audiosr; wang2025solospeech), flow-matching (wang2025flowse; liugenerative; ku2024generative), and schrödinger bridges (li2025bridgesr; jukic2024bridgese) for various restoration tasks. Despite achieving state-of-the-art results, these generative models are typically trained using likelihood-based objectives, which often misalign with human perceptual quality. Our work addresses this gap by introducing a preference-based post-training strategy applicable across these diverse generative paradigms.
@@ 2.2 Post-Training for Audio Alignment
Post-training alignment has proven effective across various generative domains, including text (ouyang2022instructgpt; rafailov2023direct; bai2022rlhf), vision (xu2023imagereward; wallace2024diffusiondpo; liu2025videoalign), speech (zhang2025advancing; sun2025f5r; hussain2025koel; tian2025preference; yao2025fine; zhang2024speechalign), music (cideron2024musicrl; lei2025levo), and audio effects (majumder2024tango2; liao2024baton). In the context of speech restoration, early efforts like MetricGAN (fu2019metricgan) used adversarial training to optimize for metrics like PESQ. More recently, preference-based methods have emerged, such as aligning with a learned NISQA predictor via PPO (kumar2025using) or using DPO with a single metric like UTMOS (li2025aligning). While single-metric alignment is effective, it risks reward hacking by ignoring that robust restoration is multi-faceted, requiring a balance of quality, fidelity, and timbre (Section 4.3). Grounded in this principle, we propose a multi-metric alignment strategy to pursue holistic improvement. We systematically show this approach yields more comprehensive gains across diverse generative paradigms while naturally mitigating reward hacking.
@@ 3.1 The GenSR-Pref Dataset: Curating a Holistic Preference Signal
Masked Generative Model (MGM). We use the pre-trained AnyEnhance (zhang2025anyenhance) model as our MGM model. It predicts acoustic tokens based on the DAC codec (kumar2024high) from partially masked sequences, effectively capturing non-autoregressive and context-aware dependencies.
Autoregressive Model (AR). We train a new model, AR+Soundstorm, which follows a two-stage pipeline common in text-to-speech (TTS) domain. It first predicts semantic tokens of the Metis codec (wang2025metis) given noisy inputs, and a second-stage non-autoregressive model similar to Soundstorm (borsos2023soundstorm) then converts these tokens into acoustic tokens for waveform synthesis.
Flow-Matching Model (FM). We train Flow-SR, a model that learns to predict the velocity field of clean mel-spectrograms from noisy inputs via optimal transport interpolation based on the Diffusion Transformer (DiT) architecture (peebles2023scalable). A pre-trained Vocos vocoder (siuzdak2023vocos) is then used to synthesize the final waveform from the generated spectrograms. Details on model architectures and training procedures can be found in Appendix B.
@@
  • Perceptual Quality (NISQA) (mittag2021nisqa): Assesses overall listening quality, including naturalness and the absence of annoying artifacts.
  • Signal-level Fidelity (DNSMOS) (reddy2022dnsmos): A composite metric evaluating signal distortion, background noise, and overall quality for a fine-grained assessment of signal integrity.
  • Content Alignment (SpeechBERTScore) (saeki2024speechbertscore): Measures semantic similarity to the ground-truth transcription to ensure linguistic content is not altered during restoration.
@@ 3.2 GenSR Model Alignment via Direct Preference Optimization
In the second stage of our approach, we align the generative speech restoration models using the GenSR-Pref dataset. We employ Direct Preference Optimization (DPO) (rafailov2023direct), a simple yet powerful technique that directly optimizes a policy to satisfy preferences without explicit reward modeling or reinforcement learning. The core idea of DPO is to re-parameterize the reward function in the standard Reinforcement Learning from Human Feedback (RLHF) objective using the optimal policy, which leads to a simple contrastive loss over preferred and dispreferred samples. In this section, we detail how we adapt and apply DPO to the three diverse generative paradigms used in our study: Autoregressive (AR), Masked Generative (MGM), and Flow-Matching (FM) models. Detailed derivations can be found in Appendix A.
@@
MGM models (chang2022maskgit; zhang2025anyenhance; wang2025metis; ju2024naturalspeech; wang2024maskgct), also known as discrete diffusion models, learn to restore a clean speech sequence $y_{0}$ from a partially masked or corrupted version $y_{t}$, where $t\in[0,T]$ is the noise step parameter that controls the amount of corruption, and $y_{T}$ is the fully masked sequence. They are trained to predict the original tokens in the masked positions, optimizing a masked language modeling objective:
@@
where $m_{t}\in\{0,1\}^{T}$ is a binary mask indicating which tokens in the sequence are masked (1) or unmasked (0). The model learns to predict the original tokens in the masked positions, effectively learning to denoise the corrupted input. DPO Derivation. Following the INTP framework (zhang2025advancing), we extend DPO to this non-autoregressive setting. The core logic remains the same, but the policy now represents the conditional probability of the full clean sequence $y_{0}$ given the corrupted version $y_{t}$. The DPO loss for MGM is thus analogous to the AR case, contrasting the log-probabilities of the preferred and dispreferred sequences, conditioned on their respective masked inputs:
@@
where $y_{t}^{w}$ and $y_{t}^{l}$ are the masked versions of the winning ($y_{0}^{w}$) and losing ($y_{0}^{l}$) sequences from our preference dataset.
@@
FM models (lipman2022flow) learn a continuous mapping from a simple prior distribution (e.g., Gaussian noise $y_{0}\sim\mathcal{N}(0,I)$) to a complex data distribution (e.g., clean speech $y_{1}$). They are trained to predict the velocity vector field $v_{\theta}(y_{t},t,x)$ that describes the probabilistic flow of data points along a linear interpolation path from the noise $y_{0}$ to the data $y_{1}$. The standard training objective minimizes the difference between the predicted velocity and the true velocity:
@@
DPO Derivation. To extend DPO to FM models, we follow recent advances in aligning diffusion-based generators (wallace2024diffusiondpo; liu2025videoalign). Instead of modeling the full data likelihood along the entire generative trajectory, we use a tractable surrogate objective that operates on a single noise step. Specifically, we approximate the DPO objective using the model’s instantaneous prediction error at a sampled timestep, treating the L2 loss between predicted and true velocities as a proxy for negative log-likelihood. The resulting loss encourages the model to reduce velocity error for preferred samples ($y_{w}$) while increasing it for dispreferred ones ($y_{l}$), yielding the following expression:
@@ 4.1 Experimental Setup
Evaluation Benchmarks and Metrics. We conduct evaluations on two major kinds of benchmarks: (1) General speech restoration (GSR) that includes denoising, dereveration, declipping and super-resolution at the same time (Voicefixer-GSR, Librivox-GSR, CCMusic-GSR), (2) Speech Enhancement (SE) under noisy and reverberant conditions (DNS-No-Reverb and DNS-With-Reverb). For a comprehensive evaluation, we not only include the original unaligned models, but also involve state-of-the-art baselines LLaSE-G1 (kang2025llase), GenSE (anonymous2025gense), MaskSR (li2024masksr), and FlowSE (wang2025flowse), and a discriminative model Voicefixer (liu2022voicefixer). Performance is measured using a comprehensive suite of objective metrics: DNSMOS (SIG, BAK, OVRL) (reddy2022dnsmos) for signal fidelity and quality, NISQA (mittag2021nisqa) for perceptual quality, SpeechBERTScore (SBERT) (saeki2024speechbertscore) for content consistency, and speaker similarity (SIM). For subjective evaluation, we conduct A/B preference tests to assess human perception. Details on the evaluation can be found in Appendix C.
@@ 4.3 Ablation Studies and Analysis
The Role of Data Source in Alignment. Finally, we explore how the source of preference data affects alignment performance. Table 4 shows the results of aligning each model with preference pairs generated from itself (in-paradigm) versus from other paradigms (cross-paradigm). This distinction is analogous to the concepts of on-policy and off-policy learning in reinforcement learning. We observe a clear trend: all three models achieve their optimal performance when aligned with their own in-paradigm preference data. To quantitatively investigate this phenomenon, we formally define the principle of in-paradigm alignment. We hypothesize that each architecture possesses a unique “preference direction”—an optimal path for improvement. To quantify this, we define a “preference vector” as the feature-space difference between winner and loser samples ($v_{\text{pref}}=f(y_{w})-f(y_{l})$) using pre-trained w2v-bert-2 model [^1] https://huggingface.co/facebook/w2v-bert-2.0 and compute their average pairwise cosine similarity. The results (Table 5) provide strong quantitative evidence for our hypothesis: the average in-paradigm similarity (diagonal values) is consistently higher than the cross-paradigm similarity (e.g., AR-AR similarity of 0.0075 vs. AR-MGM of 0.0018). This indicates that data from a model’s own paradigm provides a more consistent and targeted optimization signal.